Tim works in Nova Scotia and is paid on a monthly basis. He is receiving a...

  1. Tim works in Nova Scotia and is paid on a monthly basis. He is receiving a special pay (in addition to his 12 pays in the year) for a long service award. He is receiving an amount of $500 cash and he has not received a long service aware for 5 years. Determine which are pensionable and insurable and then calculate the CPP and EI deductions.
Answer #1

CPP stands for Canada pension plan. It is pension plan in which employees have to make contributions by taking into account TAXABLE salary or wages from employers.

In given question TIM received long service award in form CASH as it is TAXABLE that why whole $500 is pensionable amount.

EI stands for employment insurance in which taxable income limit upto $53,100 for year 2019.

So the entire $500 award is PENSIONABLE and INSURABLE as it is Cash payment and not in form of gift or awards.

Calculation for deduction for LONG SERVICE AWARD ONLY as basic pay is not known.

CPP deduction

Basic exemption $3500 for 2019.

Step 1.

Basic exemption/ 52 = $3500/52 =$67.30.

Step 2

Total PENSIONABLE income assumption $500 only as other income is not known.

$500 -$67.30 = $432.7

Step 3

Rate of CPP for 2019 is 5.10% of step 2 amount

$432.7 ×5.10 = $22.0677 for employee and matching contributions from employer as well.

Employment Insurance

Rate for EL premium is $1.62 per $100 subject to maximum taxable income of $53,100.

EL PREMIUM FOR LONG SERVICE AWARD will be $500×1.62/100=$8.1

So total contribution will be $22.0677+$8.10 =$ 30.1677.